find the sum of all 5 digit numbers that can be formed by using the digits 2,3,5,7 and 8, where each of these digits appears in each number.
Answer by Edwin McCravy(20056)


Suppose you had them all lined up in a column to add:

  23578
  23587
  23758
  23785
  23857
  23875
  25378
   ...
  85732
  87235
  87253
  87325
  87352
  87523
 +87532
=======

There are 5!=120 numbers to be added.  One 5th of
the numbers in each column are 2's, one-fifth are 3's,
and so on to one-fifth are 8's,  So since one-fifth of
120 is 24, the sum of each column is 

24*(2+3+5+7+8) = 24*(25) = 600

So there are 

600 ten-thousands +
600 thousands +
600 hundreds +
600 tens +
600 ones

or

 6000000 +
  600000 +
   60000 +
    6000 +
     600 
========
 6666600

Edwin